# Counterfactual Uncertainty Quantification of Factual Estimand of Efficacy from Before-and-After Treatment Repeated Measures Randomized Controlled Trials

### v1

@@ Abstract
The ideal estimand for comparing a new treatment $Rx$ with a control $C$ is the counterfactual efficacy $Rx:C$, the expected differential outcome between $Rx$ and $C$ if each patient were given both. While counterfactual point estimation from factual Randomized Controlled Trials (RCTs) has been available, this article shows counterfactual uncertainty quantification (CUQ), quantifying uncertainty for factual point estimates but in a counterfactual setting, is surprisingly achievable. We achieve CUQ whose variability is typically smaller than factual UQ, by creating a new statistical modeling principle called ETZ which is applicable to RCTs with Before-and-After treatment Repeated Measures, common in many therapeutic areas.
We urge caution when estimate of the unobservable true condition of a patient before treatment has measurement error, because that violation of standard regression assumption can cause attenuation in estimating treatment effects. Fortunately, we prove that, for traditional medicine in general, and for targeted therapy with efficacy defined as averaged over the population, counterfactual point estimation is unbiased. However, for targeted therapy, both Real Human and Digital Twins approaches should respect this limitation, lest predicted treatment effect in subgroups will have bias.
Keywords: Variability decomposition; Randomized Controlled Trials, Decision-making process, Confirmability.
@@ 1 Motivation and scope of application
When comparing a new treatment (which we denote by $Rx$) with a control (denoted by $C$), the ideal estimand is the counterfactual efficacy, the expected differential outcome between $Rx$ and $C$ if each patient were given both $Rx$ and $C$. Following the seminal paper Rubin (1978), causal treatment effect is defined as
@@
Most factual Randomized Controlled Trials (RCTs) are such that each patient is given either $Rx$ or $C$ but not both. However, counterfactual point estimation of counterfactual $Rx:C$ efficacy from factual studies is readily available, as random assignment of treatments to patients lets one prove that the point estimate of the differential $Rx$ and $C$ effects from modeling a factual RCT is unbiased for the counterfactual $Rx:C$ efficacy.
@@ 2 ETZ, a modeling principle
In practice, the differential effect between $Rx$ and $C$ is typically estimated from outcomes observed with some patients given $Rx$ and the others given $C$. Rubin (1978) showed that Randomized Controlled Trials (RCTs) with strongly ignorable assignment of treatments to patients allow for unbiased counterfactual point estimation of the average treatment effect from observable outcomes.
@@ 3.3 Plausibility
Intuitively, variability at the first and milestone visits and of Change may contain all the information we need to separate the $E$, $Traj$, and $Z$ variability components. Baseline measurements at the beginning of the study (visit 1) have the random intercept $Z$ and measurement error $E$ components but no random trajectory component $Traj$ since patients have not been treated yet. Change, or change-from-baseline, measurement at the milestone visit minus measurement at the first visit, has the random trajectory $Traj$ and the measurement error $E$ components (two copies of $E$ in fact) but not the random intercept component since that has been subtracted off. Measurements at the end of the study (visit $m$) have all three random components. So it seems plausible that111with the assumption variability of measurement errors is constant across the visits $Var(Y^{[1]}_{i})$, $Var(Y^{[m]}_{i})$, and $Var(Y^{[\mathrm{change}]}_{i})$ are sufficient for us to recover the the $E$, $Traj$, and $Z$ variability components.
@@ 3.5 Uncertainty reduced by counterfactual uncertainty quantification
We show below that baselining effectively replaces the patient-to-patient variability $Var(Z)$ by the measurement error variability $Var(E)$. Thus, so long as $Var(Z)>Var(E)$, baselining is beneficial. [^2] If measurement variability is larger than patient-to-patient variability, $Var(E)>Var(Z)$, then it is hopeless for that clinical trial to be informative and a different outcome measure should be chosen.
@@
and $Var(Y_{i}^{[change]})<Var(Y_{i}^{[m]})$ when $Var(E)<Var(Z)$. Figure 2 in Section 3.6 shows that baselining can reduce total variability by 73% in an example data set.
@@ 4.2 From traditional medicine to targeted therapy
A tyrosine kinase inhibitor, a small molecule with compound suffix tinib, each intending to block a specific reaction within cells, is also a targeted therapy. Example of tinibs are ibrutinib (for chronic lymphocytic leukemia CLL) and baricitinib (for rheumatoid arthritis RA).
@@ 4.3 Noisy predictor causes attenuation
Since parallel traditional medicine model is a special case of non-parallel targeted therapy model in which $\prescript{{}_{Z}}{}{\beta}^{Rx}=\prescript{{}_{Z}}{}{\beta}^{C}$, we use non-parallel targeted therapy model to illustrate attenuation in general.
@@
Take treatment $C$ as an example and similarly for $Rx$. This correlation between error and the covariate causes attenuation, because with $Z_{s}$ and $\prescript{{}_{B}}{}{E}^{[1]}_{s}$ being independent,
@@
We show below the presence of measurement error $\prescript{{}_{B}}{}{E}$ or $E$ in using baseline or any other biomarker $B$ score as a surrogate for the true patient effect $Z$ causes attenuation in estimating the regression coefficient of $B$, which in turn can cause bias in estimating $Rx:C$ efficacy. While future ETZ research might mitigate this bias, for now we show estimated counterfactual efficacy is unbiased so long as efficacy is defined as “on average”.
@@ 4.4 Possible bias caused by attenuation
In ETZ modeling, the three components $Z$, $Traj$ and $E$ follow different Normal, independent, distributions. $Z_{i}$ is generated from a Normal distribution with mean $\alpha$ and variance $Var(Z)$. $Traj^{Rx[m]}_{i}$ and $Traj^{C[m]}_{i}$ are generated from Normal distributions with means $\mu^{Rx}$ and $\mu^{C}$ and common variance $Var(Traj)$, independently. $E^{[1]}_{i}$, $E^{Rx[m]}_{i}$ and $E^{C[m]}_{i}$ are generated from Normal distributions with mean 0 and variance $Var(E)$, independently. In the counterfactual setting, outcomes $Rx$ and $C$ for patient $i$ share the same $Z_{i}$.
@@
Table 4: Variance and standard deviation (SD) of iADL of visit 1, visit $m$, and change-from-baseline of the EXPEDITION3 study. Variances of Intercept, Trajectory and Measurement Error are computed by ETZ transformation.
@@
Figure 7: The direction of bias in $Rx:C$ efficacy estimation depends on the target value of outcome $C$. Efficacy is overestimated when target $C$ value is below average, and underestimated when target $C$ value is above average. For the red points, the (horizontal) abscissa (coordinate) is outcome $C$ as predicted by $B$ and the (vertical) ordinate (coordinate) is outcome $Rx$ as predicted by $Z$. The red line is the the fitted line of $Rx$ as predicted by true $Z$ regressed on $C$ as predicted by $B$. For the teal points, both the abscissa outcome $C$ and the ordinate outcome $Rx$ are as predicted by true $Z$. The teal line is the fitted line of $Rx$ as predicted by true $Z$ regressed on $C$ as predicted by true $Z$. For easier visual perception, the $Rx$ and $C$ outcome values plotted have been partially denoised using values predicted by $Z$ after we reduce the standard deviation of $Z$ from the original 7.34 to 3.
Using baseline or, more generally, a biomarker score $B$ to predict outcome is tempting because $Z$ is not observable. However, as Figures 6 and 7 show, one should be aware that predicting expected outcome at a $B$ value off-center may incur bias if $B$ has measurement error.
@@ 4.5 Estimated average $Rx:C$ efficacy is unbiased
One intuitive way of seeing why estimated average effect is unbiased is to imagine regressing $Y$ on $Z$ in (28) versus regressing $Y$ on $B$ in (30), with a rather large number of patients. Attenuation, which occurs when $B$ is a surrogate of $Z$ but measured with error, means the slope regressing $Y$ on $B$ in (30) tends to be closer to zero, compared to the true slope regressing $Y$ on $Z$ in (28).
@@
We discuss both control-side and equipoise counterfactual estimation for targeted therapy. For control-side counterfactual estimation of targeted therapy, efficacy is measured as the difference between the factual outcome of a patient receiving $Rx$ and the corresponding counterfactual $C$ outcome:
@@
5 Going forward
ETZ is a counterfactual uncertainty quantification (CUQ) principle. This article’s default setting of $Z$ being independent of $Traj$ and the technique of extracting information from MMRM analysis are proof-of-concept in nature, verifiably plausible for traditional medicine. Since ETZ is a principle, going forward, it can be applied to other settings, employing whatever technique suitable for information extraction.
For example, response to targeted therapies for amyotrophic lateral sclerosis (ALS), typically measured on the ALS Functional Rating Scale–Revised (ALSFRS-R) scale (FDA (2021)), can be modeled with a random coefficient BAtRM model, in which case one can estimate the covariance between $Z$ and $Traj$ (e.g., using the RANDOM statement in Proc Mixed of SAS). Deriving CUQ from (18) allowing $Cov(Z,Traj)\neq 0$ will be useful future research.
Likewise, for both targeted therapy and traditional medicine, attenuation bias (31) in efficacy estimation off-center when $B$ has measurement error can potentially be mitigated by estimating the attenuation using estimated $Var(Z)$ and $Var(E)$. Working out the details will also be useful future research.
@@ 6 Key messages
ETZ is an uncertainty quantification (UQ) principle for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures (BAtRM). Conducting RCT with BAtRM is standard practice in therapeutic areas affecting more than half a billion patients worldwide. Modeling individual patient outcomes in BAtRM studies, ETZ enables counterfactual uncertainty quantification (CUQ). CUQ is typically smaller than current factual uncertainty quantification. Reduction in uncertainty quantification translates to reduction in sample size. Statistical or digital counterfactual point estimation involving a biomarker with measurement error is unbiased for efficacy in the entire population but biased in subgroups.
@@ Appendix A Glossary of terminologies
Effect of a treatment: Expected outcome given treatment $Rx$ or $C$. Efficacy between treatments: Expected differential outcome between treatments $Rx$ and $C$. Attenuation: Directional bias (within each treatment arm) in estimating slope of the treatment effect when predictor contains measurement error. Factual study: A study in which each patient is given either $Rx$ or $C$ (e.g., studies with parallel designs). Counterfactual study: A study in which each patient is given both $Rx$ and $C$ (e.g., a crossover study). Counterfactual efficacy: Expected differential between $Rx$ and $C$ outcomes of patients in a counterfactual study. Counterfactual estimation (CE): Unbiased point estimation of counterfactual efficacy from a factual study. Counterfactual uncertainty quantification (CUQ): Quantification of uncertainty (such as variance) of a counterfactual point estimate calculated as if it were from a counterfactual study. Targeted therapies: Therapies such as tyrosine-kinase inhibitors and monoclonal antibodies that target specific pathways, constituting a large portion of “personalized” or “precision” medicine. Real Human Counterfactual Estimation (RHCE): Counterfactual point estimation (CE) with counterfactual uncertainty quantification (CUQ) using each real human patient as self-control. Digital Twins Counterfactual Estimation (DTCE): Counterfactual point estimation (CE) using biomarkers to predict for each $Rx$ patient its digital $C$ twin’s outcome and vice versa, then averaging the outcome differentials.

### v3

@@ Abstract
The ideal estimand for comparing treatment $Rx$ with a control $C$ is the counterfactual efficacy $Rx:C$, the expected differential outcome between $Rx$ and $C$ if each patient were given both. One hundred years ago, Neyman (1923a) proved unbiased point estimation of counterfactual efficacy from designed factual experiments is achievable. But he left the determination of how much might the counterfactual variance of this estimate be smaller than the factual variance as an open challenge. This article shows counterfactual uncertainty quantification (CUQ), quantifying uncertainty for factual point estimates but in a counterfactual setting, is achievable for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures which are common in many therapeutic areas. We achieve CUQ whose variability is typically smaller than factual UQ by creating a new statistical modeling principle called ETZ.
We urge caution in using predictors with measurement error which violates standard regression assumption and can cause attenuation in estimating treatment effects. Fortunately, we prove that, for traditional medicine in general, and for targeted therapy with efficacy defined as averaged over the population, counterfactual point estimation is unbiased. However, for both Real Human and Digital Twins approaches, predicting treatment effect in subgroups may have attenuation bias.
Keywords: Variability decomposition, Randomized Controlled Trials, Before-and-After Repeated measures, Digital Twins, Counterfactual uncertainty quantification, ETZ modelling principle.
@@ 1 Motivation and scope of application
When comparing a new treatment (which we denote by $Rx$) with a control (denoted by $C$), the ideal estimand is the counterfactual efficacy, the expected differential outcome between $Rx$ and $C$ if each patient were given both $Rx$ and $C$. Following Rubin (1978), causal treatment effect is defined as
@@
Most factual Randomized Controlled Trials (RCTs) are such that each patient is given either $Rx$ or $C$ but not both. However, counterfactual point estimation of counterfactual $Rx:C$ efficacy from factual studies is readily available, as random assignment of treatments to patients lets one prove that the point estimate of the differential $Rx$ and $C$ effects from modeling a factual RCT is unbiased for the counterfactual $Rx:C$ efficacy. The proof that designed experiments allow for unbiased counterfactual point estimation of the average treatment effect (ATE) from observable outcomes can be traced to Neyman (1923b), which has been edited and translated as Splawa-Neyman et al. (1990). Importantly, Neyman realized that the variance of a factual point estimate in standard practice is larger than what the variance of a counterfactual point estimate would be, and considered the problem of accurately assessing the latter “worthy of future study”. See page 473 of Rubin (1990) which commented on Neyman (1923a).
@@ 2 ETZ, a modeling principle
In practice, the differential effect between $Rx$ and $C$ is typically estimated from outcomes observed with some patients given $Rx$ and others given $C$.
While it is known that randomization allows unbiased estimation of a Counterfactual estimand from factual observations, Counterfactual Uncertainty Quantification (CUQ) of such an estimate has been largely thought as requiring crossover or N-of-1 studies since Neyman (1923a). We are happy to report that, in the broadly applicable setting of BAtRM RCTs, the ETZ modeling principle enables CUQ.
@@ 3.3 Plausibility
Intuitively, variability at the first and milestone visits and of Change may contain all the information we need to separate the $E$, $Traj$, and $Z$ variability components. Baseline measurements at the beginning of the study (visit 1) have the random intercept $Z$ and measurement error $E$ components but no random trajectory component $Traj$ since patients have not been treated yet. Change, or change-from-baseline, measurement at the milestone visit minus measurement at the first visit, has the random trajectory $Traj$ and the measurement error $E$ components (two copies of $E$ in fact) but not the random intercept component since that has been subtracted off. Measurements at the end of the study (visit $m$) have all three random components. So, assuming variability of measurement errors is constant across the visits, it seems plausible that $Var(Y^{[1]}_{i})$, $Var(Y^{[m]}_{i})$, and $Var(Y^{[\mathrm{change}]}_{i})$ are sufficient for us to recover the the $E$, $Traj$, and $Z$ variability components.
@@ 3.5 Uncertainty reduced by counterfactual uncertainty quantification
We show below that baselining effectively replaces the patient-to-patient variability $Var(Z)$ by the measurement error variability $Var(E)$ and, so long as $Var(Z)>Var(E)$, baselining is beneficial.
@@
and $Var(Y_{i}^{[change]})<Var(Y_{i}^{[m]})$ when $Var(E)<Var(Z)$. (Note that if measurement variability is larger than patient-to-patient variability, $Var(E)>Var(Z)$, then it is hopeless for that clinical trial to be informative and a different outcome measure should be chosen.) Figure 2 in Section 3.6 shows that baselining can reduce total variability by 73% in an example data set.
@@ 4.2 From traditional medicine to targeted therapy
A tyrosine kinase inhibitor, a small molecule with compound suffix tinib, each intending to block a specific reaction within cells, is also a targeted therapy. Examples of tinibs are ibrutinib (for chronic lymphocytic leukemia CLL) and baricitinib (for rheumatoid arthritis RA).
@@ 4.3 Noisy predictor causes attenuation
Since the parallel traditional medicine model is a special case of non-parallel targeted therapy model in which $\prescript{{}_{Z}}{}{\beta}^{Rx}=\prescript{{}_{Z}}{}{\beta}^{C}$, we use non-parallel targeted therapy model to illustrate attenuation in general.
@@
Take treatment $C$ as an example and similarly for $Rx$. This correlation between error and the covariate causes attenuation because with $Z_{s}$ and $\prescript{{}_{B}}{}{E}^{[1]}_{s}$ being independent,
@@
We show below the presence of measurement error $\prescript{{}_{B}}{}{E}$ or $E$ in using baseline or any other biomarker $B$ score as a surrogate for the true patient effect $Z$ causes attenuation in estimating the regression coefficient of $B$, which in turn can cause bias in estimating $Rx:C$ efficacy. While future ETZ research might mitigate this bias, for now we show that estimated counterfactual efficacy is unbiased so long as efficacy is defined as “on average”.
@@ 4.4 Possible bias caused by attenuation
In ETZ modeling, the three components $Z$, $Traj$, and $E$ follow different Normal, independent, distributions. $Z_{i}$ is generated from a Normal distribution with mean $\alpha$ and variance $Var(Z)$. $Traj^{Rx[m]}_{i}$ and $Traj^{C[m]}_{i}$ are generated from Normal distributions with means $\mu^{Rx}$ and $\mu^{C}$ and common variance $Var(Traj)$, independently. $E^{[1]}_{i}$, $E^{Rx[m]}_{i}$ and $E^{C[m]}_{i}$ are generated from Normal distributions with mean 0 and variance $Var(E)$, independently. In the counterfactual setting, outcomes $Rx$ and $C$ for patient $i$ share the same $Z_{i}$.
@@
Table 4: Variance and standard deviation (SD) of iADL of visit 1, visit $m$, and change-from-baseline of the EXPEDITION3 study. Variances of Intercept, Trajectory, and Measurement Error are computed by ETZ transformation.
@@
Figure 7: The direction of bias in $Rx:C$ efficacy estimation depends on the target value of outcome $C$. Efficacy is overestimated when target $C$ value is below average, and underestimated when target $C$ value is above average. For the red points, the (horizontal) abscissa (coordinate) is outcome $C$ as predicted by $B$, and the (vertical) ordinate (coordinate) is outcome $Rx$ as predicted by $Z$. The red line is the fitted line of $Rx$ as predicted by true $Z$ regressed on $C$ as predicted by $B$. For the teal points, both the abscissa outcome $C$ and the ordinate outcome $Rx$ are as predicted by true $Z$. The teal line is the fitted line of $Rx$ as predicted by true $Z$ regressed on $C$ as predicted by true $Z$. For easier visual perception, the $Rx$ and $C$ outcome values plotted have been partially denoised using values predicted by $Z$ after we reduce the standard deviation of $Z$ from the original 7.34 to 3.
Using baseline or, more generally, a biomarker score $B$ to predict the outcome is tempting because $Z$ is not observable. However, as Figures 6 and 7 show, one should be aware that predicting the expected outcome at a $B$ value off-center may incur bias if $B$ has measurement error.
@@ 4.5 Estimated average $Rx:C$ efficacy is unbiased
One intuitive way of seeing why the estimated average effect is unbiased is to imagine regressing $Y$ on $Z$ in (28) versus regressing $Y$ on $B$ in (30), with a rather large number of patients. Attenuation, which occurs when $B$ is a surrogate of $Z$ but measured with error, means the slope regressing $Y$ on $B$ in (30) tends to be closer to zero, compared to the true slope regressing $Y$ on $Z$ in (28).
@@
We discuss both control-side and equipoise counterfactual estimations for targeted therapy. For control-side counterfactual estimation of targeted therapy, efficacy is measured as the difference between the factual outcome of a patient receiving $Rx$ and the corresponding counterfactual $C$ outcome:
@@
5 Guidelines for applying the ETZ modeling principle
ETZ is a modeling principle that enables counterfactual uncertainty quantification (CUQ) for randomized studies with Before-and-After Repeated Measures (BAtRM). Scientists faced with highly variable outcome measure (i.e., large measurement error) situations can benefit from this principle, so long as they can extract the three inputs needed by ETZ, working out details for proper CUQ in their settings themselves.
For example, response to therapies for amyotrophic lateral sclerosis (ALS) is highly variable. Outcome measured on the ALS Functional Rating Scale–Revised (ALSFRS-R) scale (FDA (2021)) can be modeled with a random coefficient BAtRM model (e.g., using the RANDOM statement in Proc Mixed of SAS), providing inputs needed by ETZ.
Scientists can apply ETZ to efficacy measured as a ratio, which is common for binary and time-to-event outcome measures, provided they are aware of the following:
  1. 1.
    Mixing not-logic-respecting efficacy measures Odds Ratio (OR) and Hazard Ratio (HR) in subgroups attenuates them, even without measurement error in (biomarker) predictors, see Liu et al. (2022);
  2. 2.
    Computer packages currently give misleading stratified analysis results for all ratio efficacy (logic-respecting or not), because the prognostic effect is ignored, see Liu et al. (2023).
However, following the Subgroup Mixable Estimation (SME) principle, correct causal inference for binary and time-to-event outcomes can be obtained for logic-respecting rato efficacy such as Relative Response (RR) and Ratio-of-Time, see Ding et al. (2016), Lin et al. (2019), the discussion paper Liu et al. (2022), and its rejoinder Liu et al. (2022).
RCTs in therapeutic areas such as immunology often are BAtRM studies with binary endpoints. For example, Morand et al. (2023) reported result of the Phase 3 study SLE-BRAVE-I, testing baricitinib for the treatment of systemic lupus erythematosus (SLE). SLE-BRAVE-I was a double-blind, randomized, placebo-controlled, parallel-group Phase 3 study. A total of 760 participants were randomly assigned to receive at least one dose of baricitinib 4 mg (n=252), baricitinib 2 mg (n=255), or placebo (n=253). Primary endpoint was the proportion of patients achieving an SLE Responder Index (SRI)-4 response at week 52 in the baricitinib 4 mg treatment group compared to the placebo group. Achieving SRI-4 is a binary outcome, a composite responder index that includes improvement in disease activity (defined as at least a 4-point improvement in the SLEDAI-2K score) without worsening of the overall condition (no increase of $\geq$ 0.3 points [10 mm] in the Physician Global Assessment from baseline) or the development of significant disease activity in new organ systems (no new BILAG A or more than one new BILAG B).
To develop treatments for diseases such as ALS (which affects an estimated 31,000 patients in the U.S.) and SLE (which affects approximately 200,000 patients in the U.S.), variability in the outcome measure is a known challenge. The CUQ principle for reducing uncertainty quantification is developed to help meet this challenge, for the greater good.
@@ 6 Key messages
ETZ is an uncertainty quantification (UQ) principle for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures (BAtRM). Conducting RCT with BAtRM is standard practice in therapeutic areas affecting more than half a billion patients worldwide. By modeling individual patient outcomes in BAtRM studies, ETZ enables counterfactual uncertainty quantification (CUQ). CUQ is typically smaller than current factual uncertainty quantification. Reduction in uncertainty quantification translates to reduction in sample size. Statistical or digital counterfactual point estimation involving a biomarker with measurement error is unbiased for efficacy in the entire population but biased in subgroups.
@@ 7 Acknowledgement
8 Conflicts of interest
X.W., Y.H., S-Y.T. declare no competing interests.
Y. L. is an employee and shareholder of Eli Lilly.
J.C.H. is a professor emeritus of the Ohio State University and a consultant to Eli Lilly.
@@ Appendix A Glossary of terminologies
BAtRM: Before-and-After treatment Repeated Measures. Effect of a treatment: Expected outcome given treatment $Rx$ or $C$. Efficacy: Expected differential outcome between treatments $Rx$ and $C$. Attenuation: Directional bias (within each treatment arm) in estimating slope of the treatment effect when predictor contains measurement error. Factual study: A study in which each patient is given either $Rx$ or $C$. Counterfactual study: A study in which each patient is given both $Rx$ and $C$. Counterfactual efficacy: Expected differential between $Rx$ and $C$ outcomes of patients in a counterfactual study. Counterfactual estimation (CE): Unbiased point estimation of counterfactual efficacy from a factual study. Counterfactual uncertainty quantification (CUQ): Quantification of uncertainty (such as variance) of a counterfactual point estimate calculated as if it were from a counterfactual study. Targeted therapies: Therapies such as tyrosine-kinase inhibitors and monoclonal antibodies that target specific pathways, constituting a large portion of “personalized” or “precision” medicine. Real Human Counterfactual Estimation (RHCE): Counterfactual point estimation (CE) with counterfactual uncertainty quantification (CUQ) using each real human patient as self-control. Digital Twins Counterfactual Estimation (DTCE): Counterfactual point estimation (CE) using biomarkers to predict for each $Rx$ patient its digital $C$ twin’s outcome and vice versa, then averaging the outcome differentials.